# Strong order-one convergence of the Euler method for random ordinary differential equations driven by semi-martingale noises

### v7

@@ 3. Bound on the pathwise approximations
We also need estimates for the Euler approximation Equation 1.2 analogous to Equation 2.3 and Equation 2.8. Notice that
@@
for $j=0,\ldots,N.$ The right hand side above depends on the time mesh. A straighforward estimate independent of the time mesh is obtained from Equation 3.1 and the fact that semi-martingales are pathwise locally bounded, almost surely:
@@
Under the 2.1, assume moreover that Equation 2.4, Equation 2.5 and Equation 2.6 hold. Then, the Euler approximation Equation 1.2 for a pathwise solution of the random ordinary differential equation Equation 1.1 satisfies
@@ 4. Semi-martingale noises
For the theory of semi-martingales, we refer the reader to [32] and [28]. The main idea is that semi-martingales are stochastic processes $\{Y_{t}\}_{t\geq 0}$ which are càdlàg, adapted, and “good integrators,” in the sense that the sums
$I_{Y}(H)=H_{0}Y_{0}+\sum_{i=1}^{n}H_{i}(Y_{t_{i+1}}-Y_{t_{i}})$
@@
$H_{t}=H_{0}\chi_{0}(t)+\sum_{i=1}^{n}H_{i}\chi_{(T_{i},T_{i+1}]},$
with the topology of uniform convergence in $(t,\omega)\in[0,\infty)\times\Omega,$ where $0=T_{1}<\ldots<T_{n},$ are stopping times, $H_{i}\in\mathcal{F}_{T_{i}},$ and $|H_{i}|<\infty,$ almost surely, to the space of finitely-valued random variables, endowed with the topology of convergence in probability. We recall the definitions and some facts about semi-martingale processes.
A function defined on a real interval is called càdlàg when it is right continuous and with left limits (“continue à droite, limite à gauche”, in French). A càdlàg process is a stochastic process $\{Y_{t}\}_{t\geq 0}$ whose sample paths are almost surely càdlàg paths. Thus, $\{Y_{t}\}_{t\geq 0}$ is càdlàg process when for almost every $\omega\in\Omega$ and for all $t\geq 0$, we have that $\lim_{s\rightarrow t^{+}}Y_{s}(\omega)=Y_{t}(\omega)$ (right continuous) and the limit $\lim_{s\rightarrow t^{-}}Y_{s}(\omega)$ exists (left limit).
@@
Almost surely, the sample paths of the jump process are zero except possibly at a countable number of points, denoted $J_{Y}$, i.e. for almost every $\omega\in\Omega$, there exists a countable set $J_{Y}(\omega)\subset I$ for which $\Delta Y_{t}(\omega)$ is zero on each connected component of $I\setminus J_{Y}(\omega)$. In fact, for any given $\varepsilon>0$, the set of times for which the jump $|\Delta Y_{t}|$ is larger than or equal to $\varepsilon$ is finite. As a consequence, each sample path is continuous except on the countable null set $J_{Y}.$
Another important process associated with a semi-martingale is the quadratic variation $\{[Y,Y]_{t}\}_{t\geq 0}$ defined as the limit of the summations of quadratic increments as the mesh is refined,
$[Y,Y]_{t}=\lim_{\begin{subarray}{c}0\leq t_{0}<\ldots<t_{n}\leq t,\;n\in\mathbb{N}\\ \max_{i=1,\ldots,n}|t_{i}-t_{i-1}|\rightarrow 0\end{subarray}}\sum_{i=1}^{n}(Y_{t_{i}}-Y_{t_{i-1}})^{2},$
or, equivalently, by the formula (which is a particular case of an “integration by parts” formula for the quadratic covariation process between two semi-martingales)
$[Y,Y]_{t}=Y_{t}^{2}-2\int_{0}^{t}Y_{s^{-}}\;\mathrm{d}Y_{s}.$
@@
The quadratic variation is a càdlàg, increasing, and adapted process [32, Theorem II.22]. The continuous part $\{[Y,Y]_{t}^{c}\}_{t\geq 0}$ of the quadratic variation is defined by
$[Y,Y]_{t}^{c}=[Y,Y]_{t}-\sum_{0\leq s\leq t}\left(\Delta Y_{s}\right)^{2},$
@@
which itself is a semi-martingale of finite-variation type. Notice that $[Y,Y]_{0}=Y_{0}^{2},$ while, for the continuous part, $[Y,Y]_{t}^{0}=[Y,Y]_{0}-\left(\Delta Y_{0}\right)^{2}=Y_{0}^{2}-Y_{0}^{2}=0.$
@@
The Bichteler-Dellacherie Theorem [32, Theorem III.47] yields a characterization of a semi-martingale as a càdlàg process $\{Y_{t}\}_{t\geq 0}$ decomposable as a sum of a finite-variation (FV) process and a local martingale process, which we can write more precisely as $Y_{t}=Y_{0}+F_{t}+Z_{t}$, with an FV process $\{F_{t}\}_{t\geq 0}$ and a local martingale process $\{Z_{t}\}_{t\geq 0}$ with $F_{0}=Z_{0}=0.$
@@
A Finite-Variation process (FV process) is a càdlàg process $\{F_{t}\}_{t\geq 0}$ whose sample paths are almost surely of finite variation, $V(\{F_{t}\}_{t};K)$, on compact subsets $K\subset[0,\infty),$ i.e.
$V(\{F_{t}\}_{t};K)=\sup_{t_{0}<\ldots<t_{n}\in K,n\in\mathbb{N}}\sum_{i=1}^{n}\|F_{t_{i}}-F_{t_{i-1}}\|<\infty.$
@@
The class of FV processes include the compound Poisson process, renewal-reward processes, inhomogeneous Poisson process, Hawkes self-exciting process, and many more. It also includes transport processes $Y_{t}=g(t,Y_{0})$ where $Y_{0}$ is a random variable and $t\mapsto g(t,y_{0})$ is a càdlàg function of finite variation, for every sample value $y_{0}=Y_{0}(\omega)$.
@@
A local martingale is an adapted, càdlàg process $\{Z_{t}\}_{t\geq 0}$ such that there exists a sequence of increasing stopping times with $T_{n}\rightarrow\infty$ almost surely, for which $\{Z_{t\wedge T_{n}}\}_{t\geq 0}$ is a uniformly integrable martingale, for each $n\in\mathbb{N},$ i.e.
$\sup_{t\geq 0}\int_{T_{n}>0}|Z_{t\wedge T_{n}}|\;\mathrm{d}\mathbb{P}<\infty.$
@@
The class of local martingales includes time-changed Brownian motions and Itô diffusion processes. For more on Itô diffusion, see also [31].
@@
With the decomposition of a semi-martingale in the form $Y_{t}=Y_{0}+F_{t}+Z_{t}$, for an FV process $\{F_{t}\}_{t\geq 0}$ and a local martingale process $\{Z_{t}\}_{t\geq 0}$ with $F_{0}=Z_{0}=0,$ we recall from [28, Lemma 18.7] that
$[Y,Y]_{t}=[Z,Z]_{t}+\sum_{0\leq s\leq t}\|\Delta F_{s}\|^{2}.$
Each step is bounded by the variation of $\{F_{t}\}_{t\geq 0},$ i.e.
@@
$\sum_{s\in J_{Y}}\|\Delta F_{t}\|\leq V(\{F_{t}\}_{t};I)<\infty.$ (4.1)
In the more general case of semi-martingale, the process $\{Y_{t}\}_{t\geq 0}$ has finite quadratic variation [32, Section II.6] on the bounded interval $I=[0,T],$ and we find, using that $[Y,Y]_{0}=Y_{0}^{2}$ and $\Delta[Y,Y]_{t}=(\Delta Y)_{t}^{2}$ componentwise (see [32, Theorem II.22]),
$\sum_{s\in J_{Y}}\|\Delta Y_{t}\|^{2}\leq\|[Y,Y]_{T}\|<\infty.$ (4.2)
Since we are considering a bounded interval $I=[0,T],$ the local martingale $Z_{t}=Z_{t\land T},$ $0\leq t\leq T,$ where $t\land T=\min\{t,T\},$ is a square-integrable martingale, and the Doob Maximal inequality (see e.g. [32, Theorem I.20]) yields
@@
which is finite by the definition of martingale. Since $F_{0}=0,$ the finite variation part is bounded by
$\sup_{t\in I}\|F_{t}\|\leq V(\{F_{t}\}_{t};I),$
@@
$\mathbb{E}\left[\sup_{t\in I}\|F_{t}\|\right]\leq\mathbb{E}\left[V(\{F_{t}\}_{t};I)\right],\qquad\mathbb{E}\left[\sup_{t\in I}\|F_{t}\|^{2}\right]\leq\mathbb{E}\left[V(\{F_{t}\}_{t};I)^{2}\right].$
@@
$\mathbb{E}\left[\sup_{t\in I}\|Y_{t}\|\right]\leq\mathbb{E}[\|Y_{0}\|]+\mathbb{E}\left[V(\{F_{t}\}_{t};I)\right]+2\mathbb{E}\left[\|Z_{T}\|^{2}\right]^{1/2}.$ (4.4)
@@
$\mathbb{E}\left[\sup_{t\in I}\|Y_{t}\|^{2}\right]\leq 3\mathbb{E}[\|Y_{0}\|^{2}]+3\mathbb{E}\left[V(\{F_{t}\}_{t};I)^{2}\right]+12\mathbb{E}[\|Z_{T}\|^{2}].$ (4.5)
For the quadratic variation part of a local martingale $\{Z_{t}\}_{t\in I},$ the Burkholder-David-Gundy inequalities [32, Theorem IV.48] relate the moments of the maximum of the local martingale with moments of its quadratic variation,
$\frac{1}{c_{p}}\mathbb{E}\left[[Z,Z]_{t}^{p/2}\right]\leq\mathbb{E}\left[\max_{0\leq s\leq t}\|Z_{t}\|^{p}\right]\leq C_{p}\mathbb{E}\left[[Z,Z]_{t}^{p/2}\right],$
for $p\geq 1$ and suitable constants $c_{p},C_{p}>0.$ Thus, we obtain the bound
$\mathbb{E}\left[[Z,Z]_{t}\right]\leq 4c_{p}\mathbb{E}\left[\|Z_{T}\|^{2}\right]<\infty,$
@@
$\mathbb{E}\left[[Y,Y]_{t}\right]\leq 4c_{2}\mathbb{E}\left[\|Z_{T}\|^{2}\right]+\mathbb{E}\left[\|V(\{F_{t}\}_{t};I)\|^{2}\right].$ (4.6)
@@
$\mathbb{E}\left[\sum_{s\in J_{Y}}\|\Delta Y_{t}\|^{2}\right]\leq 4c_{2}\mathbb{E}\left[\|Z_{T}\|^{2}\right]+\mathbb{E}\left[\|V(\{F_{t}\}_{t};I)\|^{2}\right].$ (4.7)
A fundamental tool for integration and differentiation is the chain rule. In the case of an FV process, given a continuously differentiable function $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{k}$, the following change of variables formula holds (see [32, Theorems II.31 and II.33]):
@@
The formula above is written in vectorial form, with $f$ and $Y_{t}$ of the form $f(y)=(f_{i}(y_{1},\ldots,y_{k}))_{i=1,\ldots,k}$ and $Y_{t}=((Y_{t})_{i})_{i=1,\ldots,k}$, and
$Df(y)\Delta Y_{t}=\left(\nabla f_{i}(y)\Delta Y_{t}\right)_{i=1,\ldots,k}=\left(\sum_{j=1,\ldots,k}\partial y_{j}f_{i}(y)(Y_{t})_{j}\right)_{i=1,\ldots,k}.$
When $f=f(t,y)$ depends also on $t$, one can apply the above formula to the extended process $\{(t,Y_{t})\}_{t\geq 0}$. The first component is continuous hence has no jumps and the corresponding summation term vanishes. This leads to the following change of variables formula for FV processes:
@@
In the more general case of a semi-martingale noise $\{Y_{t}\}_{t\geq 0}$, the change of variables formula Equation 4.9 takes the form (see [32, Theorems II.32 and II.33])
@@
$\quad+\frac{1}{2}\int_{0}^{t}D_{yy}f(s,Y_{s^{-}})\;\mathrm{d}[Y,Y]_{s}^{c}.$
Since $[Y,Y]_{t}^{c}$ is of finite variation, the last integral is a Lebesgue-Stieltjes integral.

### v8

@@ 3. Bound on the pathwise approximations
We also need bounds for the Euler approximation Equation 1.2 analogous to Equation 2.3 and Equation 2.8. Notice that
@@
for $j=0,\ldots,N.$ The right hand side above depends on the time mesh. A straighforward bound independent of the time mesh is obtained from Equation 3.1 and the fact that semi-martingales are pathwise locally bounded, almost surely:
@@
Under the 2.1, assume that Equation 2.4, Equation 2.5 and Equation 2.6 hold. Then, the Euler approximation Equation 1.2 for a pathwise solution of the random ordinary differential equation Equation 1.1 satisfies
@@ 4. Semi-martingale noises
For the theory of semi-martingales, we refer the reader to [1, 22, 29, 37, 41]. Starting with the approach of [41], the main idea is that semi-martingales are stochastic processes $\{Y_{t}\}_{t\in I}$ which are càdlàg, adapted, and “good integrators”, in the sense that the sums
$I_{Y}(H)=H_{0}Y_{0}+\sum_{i=1}^{n}H_{i}(Y_{t_{i}}-Y_{t_{i-1}})$
@@
$H_{t}=H_{0}\chi_{0}(t)+\sum_{i=1}^{n}H_{j}\chi_{(T_{i-1},T_{i}]},$
to the space of a.s. finite-valued random variables, where $0=T_{0}<T_{1}<\ldots<T_{n},$ are stopping times, $H_{j}\in\mathcal{F}_{T_{j}},$ and $|H_{j}|<\infty$ almost surely, with the former space endowed with the topology of uniform convergence in $(t,\omega)\in[0,\infty)\times\Omega,$ and the latter, with the topology of convergence in probability. We recall the definitions and some facts about (scalar or vector-valued) semi-martingale processes.
A function defined on a real interval is called càdlàg when it is right continuous and with left limits (“continue à droite, limite à gauche”, in French). A càdlàg process is a stochastic process $\{Y_{t}\}_{t\in I}$ whose sample paths are almost surely càdlàg paths. Thus, $\{Y_{t}\}_{t\in I}$ is a càdlàg process when for almost every $\omega\in\Omega$ and for all $t\in I$, we have that $\lim_{s\rightarrow t^{+}}Y_{s}(\omega)=Y_{t}(\omega)$ (right continuous) and the limit $\lim_{s\rightarrow t^{-}}Y_{s}(\omega)$ exists (left limit).
@@
Almost surely, the sample paths of the jump process are zero except possibly at a countable number of points, denoted $J_{Y}$, i.e. for almost every $\omega\in\Omega$, there exists a countable set $J_{Y}(\omega)\subset I$ for which $\Delta Y_{t}(\omega)$ is zero on each connected component of $I\setminus J_{Y}(\omega)$. In fact, for any given $\varepsilon>0$, the set of times for which the jump $\|\Delta Y_{t}\|$ is larger than or equal to $\varepsilon$ is finite. As a consequence, each sample path is continuous except on the countable null set $J_{Y}.$
Another important process associated with a semi-martingale is the quadratic variation $\{[Y,Y]_{t}\}_{t\in I}$. It is defined either as a suitable limit of the summations of quadratic steps $(Y_{T_{j}^{n}}-Y_{T_{j-1}^{n}})^{2}$ over random partitions made of stopping times $T_{j}^{n}$ or, equivalently, as a suitable integral [41, Section II.6]. In the vector-valued case, denoting $Y_{t}=(Y_{t}^{(l)})_{l=1,\ldots,k}$, with the vectorial components $Y_{t}^{(l)},$ the quadratic process becomes matrix-valued, i.e. $Y,Y]_{t}=([Y^{(l_{1})},Y^{(l_{2})}]_{t})_{l_{1},l_{2}=1,\ldots,k},$ with the appearance of the quadratic variations and covariations. The equivalent integral definition takes, more precisely, the form of an “integration by parts” formula for two semi-martingales,
$Y^{(l_{1})},Y^{(l_{2})}]_{t}=Y_{t}^{(l_{1})}Y_{t}^{(l_{2})}-\int_{0}^{t}Y_{s^{-}}^{(l_{1})}\;\mathrm{d}Y_{s}^{(l_{2})}-\int_{0}^{t}Y_{s^{-}}^{(l_{2})}\;\mathrm{d}Y_{s}^{(l_{1})}.$
@@
The quadratic variation is a càdlàg, increasing, and adapted process [41, Theorem II.22]. The continuous part $\{[Y,Y]_{t}^{c}\}_{t\in I}$ of the quadratic variation is defined by
$Y,Y]_{t}^{c}=[Y,Y]_{t}-\sum_{0\leq s\leq t}\Delta Y_{s}\Delta Y_{s}^{\operatorname{tr}},$
@@
which itself is a semi-martingale of finite-variation type, and where $Y_{t}^{\operatorname{tr}}$ is the transpose of $Y_{t}.$ Notice that $Y,Y]_{0}=Y_{0}^{2},$ while, for the continuous part, $Y,Y]_{0}^{c}=[Y,Y]_{0}-\Delta Y_{0}\Delta Y_{0}^{\operatorname{tr}}=Y_{0}Y_{0}^{\operatorname{tr}}-Y_{0}Y_{0}^{\operatorname{tr}}=0.$
@@
The Bichteler-Dellacherie Theorem [41, Theorem III.47] yields a characterization of a semi-martingale as a càdlàg process $\{Y_{t}\}_{t\in I}$ decomposable as a sum of a finite-variation (FV) process and a local martingale process, which we can write in standard form as $Y_{t}=Y_{0}+F_{t}+Z_{t}$, with an FV process $\{F_{t}\}_{t\in I}$ and a local martingale process $\{Z_{t}\}_{t\in I}$ with $F_{0}=Z_{0}=0.$
@@
A Finite-Variation process (FV process) is a càdlàg process $\{F_{t}\}_{t\in I}$ whose sample paths are almost surely of finite variation $V({\{F_{t}\}_{t\in I}};K)$ on compact subsets $K\subset I,$ i.e.
$V(\{F_{t}\}_{t\in I};K)=\sup_{t_{0}<\ldots<t_{n}\in K,n\in\mathbb{N}}\sum_{j=1}^{n}\|F_{t_{j}}-F_{t_{j-1}}\|<\infty.$
This total variation induces a positive measure $\|\mathrm{d}F_{t}\|,$ which is the infinitesimal of this total variation, so that $\|\mathrm{d}F_{t}\|([a,b])=V({\{F_{t}\}_{t\in I}};[a,b]),$ for every $a,b]\subset I.$
@@
The class of FV processes includes the compound Poisson process, renewal-reward processes, inhomogeneous Poisson process, Hawkes self-exciting process, and many more. It also includes transport processes $Y_{t}=g(t,Y_{0})$ where $Y_{0}$ is a random variable and $t\mapsto g(t,y_{0})$ is a càdlàg function of finite variation, for every sample value $y_{0}=Y_{0}(\omega)$.
@@
A local martingale is an adapted, càdlàg process $\{Z_{t}\}_{t\in I}$ such that there exists a sequence of increasing stopping times $T_{n}$ with $T_{n}\rightarrow\infty$ almost surely, for which $\{Z_{t\wedge T_{n}}1_{T_{n}>0}\}_{t\in I}$ is a uniformly integrable martingale, for each $n\in\mathbb{N}.$ Here, $t\land T=\min\{t,T\},$ the function $1_{T_{n}>0}$ is the indicator function of the set $\{\omega;\;T_{n}(\omega)>0\},$ and uniform integrability means
$\lim_{k\rightarrow\infty}\sup_{t\in I}\int_{\{T_{n}>0\}\cap\{\|Z_{t\wedge T_{n}}\|\geq k\}}\|Z_{t\wedge T_{n}}\|\;\mathrm{d}\mathbb{P}=0.$
@@
The class of local martingales includes time-changed Brownian motions and Itô diffusion processes. For more information on Itô diffusion, see also [40].
@@
With the decomposition $Y_{t}=Y_{0}+F_{t}+Z_{t}$ described above, we recall from [37, Lemma 18.7] that
$Y,Y]_{t}=[Z,Z]_{t}+\sum_{0\leq s\leq t}\Delta F_{s}\Delta F_{s}^{\operatorname{tr}}.$
Each step is bounded by the variation of $\{F_{t}\}_{t\in I},$ i.e.
@@
$\sum_{s\in J_{Y}}\|\Delta F_{s}\|\leq V(\{F_{t}\}_{t\in I};I)<\infty.$ (4.1)
In the more general case of a semi-martingale, the process $\{Y_{t}\}_{t\in I}$ has finite quadratic variation [41, Section II.6] on the bounded interval $I=[0,T],$ and we find, using that $Y,Y]_{0}=Y_{0}Y_{0}^{\operatorname{tr}}$ and $\Delta[Y,Y]_{t}=\Delta Y_{t}\Delta Y_{t}^{\operatorname{tr}}$ (see [41, Theorem II.22]), that
$\sum_{s\in J_{Y}}\|\Delta Y_{s}\|^{2}\leq\|[Y,Y]_{T}\|<\infty.$ (4.2)
By Doob’s maximal quadratic inequality (see e.g. [41, Theorem I.20]), if $Z_{t}=Z_{t\land T},$ $0\leq t\leq T,$ is a square-integrable martingale, then
@@
Since $F_{0}=0,$ the finite variation part is bounded by
$\sup_{t\in I}\|F_{t}\|\leq V(\{F_{t}\}_{t\in I};I),$
@@
$\mathbb{E}\left[\sup_{t\in I}\|F_{t}\|\right]\leq\mathbb{E}\left[V(\{F_{t}\}_{t\in I};I)\right],\qquad\mathbb{E}\left[\sup_{t\in I}\|F_{t}\|^{2}\right]\leq\mathbb{E}\left[V(\{F_{t}\}_{t\in I};I)^{2}\right].$
@@
$\mathbb{E}\left[\sup_{t\in I}\|Y_{t}\|\right]\leq\mathbb{E}[\|Y_{0}\|]+\mathbb{E}\left[V(\{F_{t}\}_{t\in I};I)\right]+2\mathbb{E}\left[\|Z_{T}\|^{2}\right]^{1/2},$ (4.4)
@@
$\mathbb{E}\left[\sup_{t\in I}\|Y_{t}\|^{2}\right]\leq 3\mathbb{E}[\|Y_{0}\|^{2}]+3\mathbb{E}\left[V(\{F_{t}\}_{t\in I};I)^{2}\right]+12\mathbb{E}[\|Z_{T}\|^{2}].$ (4.5)
For the quadratic variation part of a local martingale $\{Z_{t}\}_{t\in I},$ the Burkholder-David-Gundy inequalities (see e.g. [41, Theorem IV.48] and [35]) relate the moments of the maximum of the local martingale with moments of its quadratic variation,
$\frac{1}{\gamma_{p}}\mathbb{E}\left[\|[Z,Z]_{t}\|^{p/2}\right]\leq\mathbb{E}\left[\max_{0\leq s\leq t}\|Z_{s}\|^{p}\right]\leq\tilde{\gamma}_{p}\mathbb{E}\left[\|[Z,Z]_{t}\|^{p/2}\right],$
for $p\geq 1$ and suitable constants $\gamma_{p},\tilde{\gamma}_{p}>0.$ Thus, we obtain the bound
$\mathbb{E}\left[\|[Z,Z]_{t}\|\right]\leq 4\gamma_{2}\mathbb{E}\left[\|Z_{T}\|^{2}\right]<\infty,$ (4.6)
@@
$\mathbb{E}\left[\|[Y,Y]_{t}\|\right]\leq 4\gamma_{2}\mathbb{E}\left[\|Z_{T}\|^{2}\right]+\mathbb{E}\left[\|V(\{F_{t}\}_{t\in I};I)\|^{2}\right].$ (4.7)
@@
$\mathbb{E}\left[\sum_{s\in J_{Y}}\|\Delta Y_{s}\|^{2}\right]\leq 4\gamma_{2}\mathbb{E}\left[\|Z_{T}\|^{2}\right]+\mathbb{E}\left[\|V(\{F_{t}\}_{t\in I};I)\|^{2}\right].$ (4.8)
The polarization identity
$Y^{(l_{1})},Y^{(l_{2})}]=\frac{1}{2}\left([Y^{(l_{1})}+Y^{(l_{2})},Y^{(l_{1})}+Y^{(l_{2})}]-[Y^{(l_{1})},Y^{(l_{1})}]-[Y^{(l_{2})},Y^{(l_{2})}]\right)$
implies that each quadratic covariation is a difference of quadratic variations, which are increasing processes, so that each quadratic covariation is also of finite variation (see [41, Corollary 1 of Theorem II.22]). In particular, the continuous part of the quadratic local martingale term $\{Z_{t}\}_{t\in I}$ is of finite variation and it follows from the definition of the continuous part of the quadratic variation, Equation 4.7, and Equation 4.6 that
$\mathbb{E}\left[V(\{[Z,Z]_{t}^{c}\}_{t\in I};I)\right]\leq 5\gamma_{2}\mathbb{E}\left[\|Z_{T}\|^{2}\right]<\infty.$ (4.9)
A fundamental tool for integration and differentiation is the chain rule. In the case of an FV process, given a continuously differentiable function $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$, the following change of variables formula holds (see [41, Theorems II.31 and II.33]):
@@
The formula above is written in vectorial form, with $f$ and $Y_{t}$ of the form $f(y)=(f_{i}(y_{1},\ldots,y_{k}))_{i=1,\ldots,d}$ and $Y_{t}=(Y_{t}^{(l)})_{l=1,\ldots,k}$, and
$Df(y)\Delta Y_{t}=\left(\nabla f_{i}(y)\Delta Y_{t}\right)_{i=1,\ldots,d}=\left(\sum_{l=1,\ldots,k}\partial y_{l}f_{i}(y)\Delta Y_{t}^{(l)}\right)_{i=1,\ldots,d}.$
When $f=f(t,y)$ depends also on $t$, one can apply the above formula to the extended process $\{(t,Y_{t})\}_{t\in I}$. The first component is continuous, hence it has no jumps, and the corresponding summation term vanishes. This leads to the following change of variables formula for FV processes:
@@
In the more general case of a semi-martingale noise $\{Y_{t}\}_{t\in I}$, the change of variables formula Equation 4.11 takes the form (see e.g. [41, Theorems II.32 and II.33], [1, Section 1.3.3], [29, Theorem 3.6], or [22, Theorem 9.35])
@@
$\quad+\frac{1}{2}\int_{0}^{t}D_{yy}f(s,Y_{s^{-}})\;\mathrm{d}[Y,Y]_{s}^{c},$
where each component of $D_{yy}f(s,Y_{s^{-}})\;\mathrm{d}[Y,Y]_{s}^{c}$ is given by
$\left(D_{yy}f(s,Y_{s^{-}})\;\mathrm{d}[Y,Y]_{s}^{c}\right)^{(i)}=\sum_{l_{1},l_{2}=1,\ldots,d}\partial_{y_{l_{1}}}\partial_{y_{l_{2}}}f_{i}(s,Y_{s^{-}})\;\mathrm{d}[Y^{(l_{1})},Y^{(l_{2})}]_{s}^{c}.$
Since $Y,Y]_{t}^{c}$ is of finite variation, the last integral in Equation 4.12 is a Lebesgue-Stieltjes integral.
@@
Finally, a fundamental identity is the Itô isometry, which, in the case of a vector-valued local martingale $\{Z_{t}\}_{t\in I}$ and an adapted matrix-valued process $\{A_{t}\}_{t\in I},$ takes the form (e.g. apply [29, Chapter 3, Proposition 2.17] to the expanded coordinate terms)
$\mathbb{E}\left[\left\|\int_{0}^{t}A_{s}\;\mathrm{d}Z_{s}\right\|^{2}\right]=\mathbb{E}\left[\int_{0}^{t}A_{s}^{\operatorname{tr}}A_{s}\odot\mathrm{d}[Z,Z]_{s}\right]\\ =\mathbb{E}\left[\int_{0}^{t}\sum_{l_{1},l_{2}=1}^{k}\sum_{i=1}^{d}A_{s}^{(il_{1})}A_{s}^{(il_{2})}\;\mathrm{d}[Z^{(l_{1})},Z^{(l_{2})}]_{s}\right],$ (4.13)
provided each component $\{A_{t}^{il}\}_{t\in I}$ is square integrable for the integration with respect to $\{[Z^{(l)},Z^{(l)}]_{t}\}_{t\in I}$, and where $\odot$ denotes the Hadamard product of two matrices of same dimensions. Recall that the steps of the cross variation matrix $Z,Z]_{s}$ are positive semi-definite, so the right hand side is nonnegative.